if the third term of gp is two divided by nine and sum of infinite terms is three divided by two then find the series

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
ar2=2/9
a/(1-r)=3/2
so division will lead to 
r2(1-r) =2/9*2/3
r2(1-r)=4/27
r=2/3
So the a=1/2 and the seris be ½ , 1/3 , 2/9, 4/27 ….